A ray of light in corn oil is incident striking with an angle of incidence of 40 degrees on water. Calculate the angle of refraction to the nearest degree


1
Expert's answer
2021-07-09T08:32:17-0400

The angle of refraction is given by the Snell's law:


"\sin i\cdot n_1 = \sin r\cdot n_2"

where "i = 40\degree" is the angle of incidence, "r" is the angle of refraction, and "n_1 = 1.47, n_2 = 1.33" are the indices of refraction of the oil and water respectively.

Thus, obtain:


"r = \arcsin\left( \sin i\cdot \dfrac{n_1}{n_2} \right) = \arcsin\left( \sin40\degree\cdot \dfrac{1.47}{1.33} \right) \approx 45\degree"

Answer. "45\degree".